• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name nanda

    Log entry time 01:32:51 on January12,2003

    Entry number 91533

    keyword=Compton Update

    The beam quality in the compton chicane is bad. we have 9kHz background rate in the photon calorimeter at 8uA at 4 GeV with laser off. With good beam this should be 1 to 2 kHz. The rate only goes to 17kHz with laser on. Rather poor signal to noise ratio to get a meaningful polarization number. A beam retune is necessary.

    The optical cavity is running reliably at about 1kW power.

    Frequent DAQ crashes with ROC1 and ROC2. Made access to power cycle vme crates. Seems to have solve it for now.

    Even with 2:1 signal to noise ratio we decided to keep taking data through the night.
    for analysis later.

    The compton helicity signals work fine. No bad timings.

    Stay tuned...